# Experimental demonstration of spontaneous symmetry breaking with emergent multi-qubit entanglement

### v1

@@ Abstract
Spontaneous symmetry breaking (SSB) is crucial to the occurrence of phase transitions. Once a phase transition occurs, a quantum system presents degenerate eigenstates that lack the symmetry of the Hamiltonian. After crossing the critical point, the system is essentially evolved to a quantum superposition of these eigenstates until decoherence sets in. Demonstrations of SSB with quantum-mechanical characters have been observed in superradiant phase transitions with a single qubit coupled to a single photonic mode. We here propose and demonstrate SSB in the quantum phase transition of a multi-qubit system, governed by the competition between individual drivings and intra-qubit interactions. The quantum-mechanical SSB is manifested by the emergence of a multi-qubit cat state, formed by two degenerate symmetry-breaking eigenstates. We synthesize the theoretical model in a circuit quantum electrodynamics system, where 6 Xmon qubits are coupled in an all-to-all manner through virtual photon exchange mediated by a resonator. The observed nonclassical features go beyond the conventional description of SSB, shedding new light on phase transitions for quantum many-body systems.
@@
The most typical example for illustrating SSB is a classical model, where a point-like particle moves along a straight line under the influence of a sombrero potential, given by $V(x)=\Lambda x^{4}-\mu x^{2}$ with $\Lambda,\mu>0$ [12]. The particle has two different positions of stable equilibrium, $x=\pm\sqrt{\mu/2\Lambda}$. Although the potential possesses a global spatial-inversion symmetry, i.e., $V(-x)=V(x)$, each of the two equilibrium positions loses this symmetry. Which equilibrium position the particle chooses to stay at depends on perturbation. This is the standard understanding of SSB, which, however, does not grasp the quantum-mechanical feature of SSB. For quantum systems, the superposition principle allows degenerate stable symmetry-breaking eigenstates to be superimposed with each other, forming a symmetric one. As such, SSB is said to occur if there exist two or more degenerate eigenstates, each of which lacks the symmetry of the governing Hamiltonian [12]. For an isolated quantum system initially with the same symmetry as the Hamiltonian, its symmetry will not be changed. To keep this symmetry, the system should evolve to the symmetric superposition of the corresponding degenerate eigenstates after a quantum phase transition. For a real quantum system, it is the environment-induced decoherence that destroys the quantum coherence, forcing the system to choose one of the symmetry-breaking eigenstates as the reality [13, 14, 15].
@@
We here propose and demonstrate an alternative scheme for observing SSB in the $N$-qubit Lipkin-Meshkov-Glick (LMG) model [38], where each qubit is subjected to a continuous drive and coupled to all other qubits. By slowly tuning the driving parameter, the system can evolve from an initial symmetric product state to a superposition of two degenerate symmetry-breaking eigenstates of the Hamiltonian, as a consequence of the SSB. We demonstrate this SSB process in a superconducting circuit, where 6 qubits are nearly homogeneously coupled to one another with the assistance of a bus resonator. We investigate how the symmetry is broken during the quench, and verify the quantum entanglement of the output state by measuring multi-qubit quantum correlations and mapping out the Wigner function. Our results reveal the fundamental distinction between SSB of quantum multipartite systems and that of classical ones.
@@
where $\Omega$ is the driving strength, $\lambda$ denotes the qubit-qubit coupling, $S_{x/z}=\frac{1}{2}\stackrel{{\scriptstyle N}}{{\mathrel{\mathop{\sum}\limits_{j=1}}}}\sigma_{j}^{x/z}$ is the collective angular operator, and $\sigma_{j}^{x/z}$ denotes the Pauli operators of the $j$th qubit. The Hamiltonian has $Z_{2}$ symmetry, defined as the joint parity: ${\cal P}_{x}=\otimes_{j=1}^{N}\sigma_{j}^{x}$. The system’s ground-state behavior depends on the competition of longitudinal couplings and transverse drivings. When $\Omega,\lambda>0$ and $\Omega\gg\lambda$, the longitudinal couplings are suppressed, so that the system presents a unique ground state of the form $\left|X\right\rangle=\otimes_{j=1}^{N}\left|+_{j}\right\rangle$, where |±$\left|\pm_{j}\right\rangle=\frac{1}{\sqrt{2}}(\left|g_{j}\right\rangle\pm\left|e_{j}\right\rangle)$. Here $\left|g_{j}\right\rangle$ and $\left|e_{j}\right\rangle$ denote the lower and upper levels of the $j$th qubit. This eigenstate has the same symmetry as the Hamiltonian, i.e., ${\cal P}_{x}\left|X\right\rangle=\left|X\right\rangle$. When $\Omega\ll\lambda$, the transverse driving only slightly shifts the energy levels of the eigenstates $\left|J,m\right\rangle$ ($\left|J,m\right\rangle$ is the eigenstate of $S_{z}$ with eigenvalue $m$ and total angular momentum quantum number $J$. See Supplementary Material [39] Sec. S1 for the detailed derivation). The most remarkable feature of the system in the strong-coupling regime is that it has two degenerate ground states, $\left|J,J\right\rangle=\otimes_{j=1}^{N}\left|e_{j}\right\rangle$ and $\left|J,-J\right\rangle=\otimes_{j=1}^{N}\left|g_{j}\right\rangle$, each breaking the system’s $Z_{2}$ symmetry.
@@
To characterize the state evolution, it is necessary to be able to distinguish between $\left|\psi_{0}\right\rangle=\left|+_{j}\right\rangle^{\otimes 6}$ and $\left|\psi_{1}\right\rangle=\frac{1}{\sqrt{2}}(\left|g_{j}\right\rangle^{\otimes 6}+\left|e_{j}\right\rangle^{\otimes 6})$ that are the ground states in the limits $\lambda/\Omega\rightarrow 0$ and $\Omega/\lambda\rightarrow 0$, respectively. Though both states possess a zero z-axis magnetization $\sum_{j=1}^{N}\left\langle\sigma_{j}^{z}\right\rangle$, they have very different population distributions. For $\left|\psi_{0}\right\rangle$, the probability of all qubits being populated in $\left|g\right\rangle$ is $P_{0}^{e}=1/2^{6}$, while that for $\left|\psi_{1}\right\rangle$ is $1/2$. The measured population, $P_{0}^{e}$, as a function of the quench time, is displayed in Fig. 2(b). As expected, $P_{0}^{e}$ is increased monotonously with the quench process, during which $\Omega/\lambda$ is decreased. To reveal the system evolution more detailedly, we display the population distributions measured at $\tau=0$, $\tau=50$, $\tau=100$, and $\tau=150$ ns in Figs. 2(c), 2(d), 2(e) and 2(f), respetively, which confirm the system population undergoes a significant change during the quench, transformed from a binomial distribution to a pattern dominated by the two maximally distinct components $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$. Each of these components corresponds to a symmetry-breaking order with respect to the z-axis magnetization. We note that the dynamics of the qubit chain deviates from a unitary process due to dissipation.
@@
For the initial symmetric state, this longitudinal correlation is equal to 0. With the decrease of $\Omega$, this correlation grows, and reaches the maximum $1$ when the system is entirely populated by $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$, independent of their quantum coherence. Figure 3(c) displays this correlation versus the quench time $\tau$. The result illustrates the SSB process, during which the system loses the $Z_{2}$ symmetry, as a consequence of the increasing longitudinal correlations.
@@
In conclusion, we have presented a demonstration of the SSB quantum phase transition in the LMG model. By tuning the competition between the external driving and intra-qubit interaction, we can move the system from the $Z_{2}$-symmetric phase to the symmetry-breaking phase, where the two degenerate highest eigenstates of the governing Hamiltonian, superimposed with each other, form a cat state. We demonstrate the theoretical prediction with a 6-qubit LMG model in a superconducting processor, where each of the 6 qubits is effectively coupled to all others through virtual photon exchange, mediated by a bus resonator dispersively coupled to these qubits. By gradually decreasing the strength of individual drivings, we observe the evolution of the system from a product state into an entangled GHZ state which is equivalent to SSB in modern understanding. The nonclassicality of the produced state is confirmed and characterized by the quantum coherence measurement and multi-qubit Wigner tomography. Besides fundamental interest, the realized SSB phase transition offers a dephasing-insensitive strategy for large-scale entanglement manipulation.

### v2

@@ Abstract
Spontaneous symmetry breaking (SSB) is crucial to the occurrence of phase transitions. Once a phase transition occurs, a quantum system presents degenerate eigenstates that lack the symmetry of the Hamiltonian. After crossing the critical point, the system is essentially evolved to a quantum superposition of these eigenstates until decoherence sets in. Despite the fundamental importance and potential applications in quantum technologies, such quantum-mechanical SSB phenomena have not been experimentally explored in many-body systems. We here present an experimental demonstration of the SSB process in the Lipkin-Meshkov-Glick model, governed by the competition between the individual driving and intra-qubit interaction. The model is realized in a circuit quantum electrodynamics system, where 6 Xmon qubits are coupled in an all-to-all manner through virtual photon exchange mediated by a resonator. The observed nonclassical correlations among these qubits in the symmetry-breaking region go beyond the conventional description of SSB, shedding new light on phase transitions for quantum many-body systems.
@@
The most typical example for illustrating SSB is a classical model, where a point-like particle moves along a straight line under the influence of a sombrero potential, given by $V(x)=\Lambda x^{4}-\mu x^{2}$ with $\Lambda,\mu>0$ [12]. The particle has two different positions of stable equilibrium, $x=\pm\sqrt{\mu/2\Lambda}$. Although the potential possesses a global spatial-inversion symmetry, i.e., $V(-x)=V(x)$, each of the two equilibrium positions loses this symmetry. Which equilibrium position the particle chooses to stay at depends on perturbation. This is the standard understanding of SSB, which, however, does not grasp the quantum-mechanical feature of SSB. For quantum systems, the superposition principle allows degenerate stable symmetry-breaking eigenstates to be superimposed with each other, forming a symmetric one. As such, SSB is said to occur if there exist two or more degenerate eigenstates, each of which lacks the symmetry of the governing Hamiltonian [12]. When the system initially has the same symmetry as the Hamiltonian, it should evolve to the symmetric superposition of the corresponding degenerate eigenstates after a quantum phase transition under the Hamiltonian dynamics. For an open quantum system, it is the environment-induced decoherence that destroys the quantum coherence, forcing the system to choose one of the symmetry-breaking eigenstates as the reality [13, 14, 15].
@@
We here demonstrate the purely quantum-mechanical SSB, featuring emergent GHZ entanglement, in the multi-qubit Lipkin-Meshkov-Glick (LMG) model [38], governed by a Hamiltonian featuring the competition between individual continuous drivings and intra-qubit interactions. The LMG model is synthesized in a superconducting circuit, where 6 Xmon qubits are nearly homogeneously coupled to one another with the assistance of a bus resonator connected to all these qubits. By slowly decreasing the driving strength, we realize the SSB process, during which the multi-qubit system evolves from an initial symmetric product state to a superposition of two degenerate symmetry-breaking eigenstates of the governing Hamiltonian. Through the real-time probing of the average two-qubit longitudinal correlation during the quench process, we observe the progressive improvement of the $Z_{2}$ symmetry breaking degree. We verify the quantum character of the SSB by measuring multi-qubit transverse quantum correlations and mapping out the Wigner function. Our results reveal the fundamental distinction between the SSB of quantum many-body systems and that of classical ones, and offer a dephasing-insensitive strategy for multi-qubit entanglement engineering in solid-state systems.
@@
where $\Omega$ is the driving strength, $\lambda$ denotes the qubit-qubit coupling, $S_{x/z}=\frac{1}{2}\stackrel{{\scriptstyle N}}{{\mathrel{\mathop{\sum}\limits_{j=1}}}}\sigma_{j}^{x/z}$ is the collective angular operator, and $\sigma_{j}^{x/z}$ denotes the Pauli operators of the $j$th qubit. The Hamiltonian has the $Z_{2}$ symmetry, defined as the joint parity: ${\cal P}_{x}=\otimes_{j=1}^{N}\sigma_{j}^{x}$. The system’s ground-state behavior depends on the competition of longitudinal couplings and transverse drivings. When $\Omega,\lambda>0$ and $\Omega\gg\lambda$, the longitudinal couplings are suppressed, so that the system presents a unique ground state of the form $\left|X\right\rangle=\otimes_{j=1}^{N}\left|+_{j}\right\rangle$, where |±$\left|\pm_{j}\right\rangle=\frac{1}{\sqrt{2}}(\left|g_{j}\right\rangle\pm\left|e_{j}\right\rangle)$. Here $\left|g_{j}\right\rangle$ and $\left|e_{j}\right\rangle$ denote the lower and upper levels of the $j$th qubit. This eigenstate has the same symmetry as the Hamiltonian, i.e., ${\cal P}_{x}\left|X\right\rangle=\left|X\right\rangle$. When $\Omega\ll\lambda$, the transverse driving only slightly shifts the energy levels of the eigenstates $\left|J,m\right\rangle$ ($\left|J,m\right\rangle$ is the eigenstate of $S_{z}$ with eigenvalue $m$ and total angular momentum quantum number $J$. See Supplementary Material [39] Sec. S1 for the detailed derivation). The most remarkable feature of the system in the strong-coupling regime is that it has two degenerate ground states, $\left|J,J\right\rangle=\otimes_{j=1}^{N}\left|e_{j}\right\rangle$ and $\left|J,-J\right\rangle=\otimes_{j=1}^{N}\left|g_{j}\right\rangle$, each breaking the system’s $Z_{2}$ symmetry.
@@
To characterize the state evolution, it is necessary to be able to distinguish between $\left|\psi_{0}\right\rangle=\left|+_{j}\right\rangle^{\otimes 6}$ and $\left|\psi_{1}\right\rangle=\frac{1}{\sqrt{2}}(\left|g_{j}\right\rangle^{\otimes 6}+\left|e_{j}\right\rangle^{\otimes 6})$ that are the ground states in the limits $\lambda/\Omega\rightarrow 0$ and $\Omega/\lambda\rightarrow 0$, respectively. Though both states possess a zero z-axis magnetization $\sum_{j=1}^{N}\left\langle\sigma_{j}^{z}\right\rangle$, they have very different population distributions. For $\left|\psi_{0}\right\rangle$, the probability of all qubits being populated in $\left|g\right\rangle$ is $P_{0}^{e}=1/2^{6}$, while that for $\left|\psi_{1}\right\rangle$ is $1/2$. The measured population, $P_{0}^{e}$, as a function of the quench time, is displayed in Fig. 2(b). As expected, $P_{0}^{e}$ is increased monotonously with the quench process, during which $\Omega/\lambda$ is decreased. To reveal the system evolution more detailedly, we display the population distributions measured at $\tau=0$, $\tau=50$, $\tau=100$, and $\tau=150$ ns in Figs. 2(c), 2(d), 2(e) and 2(f), respetively, which confirm the system population undergoes a significant change during the quench, transformed from a binomial distribution to a pattern dominated by the two maximally distinct components $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$. Each of these components corresponds to a symmetry-breaking order with respect to the z-axis magnetization. We note that the dynamics of the system deviates from a unitary process due to dissipation.
@@
This longitudinal correlation is equal to 0 for the $Z_{2}$-symmetric eigenstate $|X\rangle$, and equals to 1 for any combination of the $Z_{2}$-symmetry-breaking eigenstates $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$, independent of their quantum coherence. This implies that ${\cal C}_{2}^{L}$ can be used as a measure to characterize the SSB degree. For the ideal LMG model, ${\cal C}_{2}^{L}$ would monotonously increase when $\Omega$ is adiabatically decreased, tending to 1 when $\Omega/\lambda\to 0$. Figure 3(c) displays ${\cal C}_{2}^{L}$ measured for different quench times. As expected, with the decrease of $\Omega$, it is progressively increased until $\tau=150$ ns. Due to the experimental imperfections, the maximum of ${\cal C}_{2}^{L}$ is limited to 0.67. The populations of $\left|g_{j}\right\rangle^{\otimes 6}$ and $\left|e_{j}\right\rangle^{\otimes 6}$ corresponding to this maximum are 0.34 and 0.32, respectively.
@@
In conclusion, we have presented a demonstration of the SSB in a multi-qubit LMG model, engineered with a superconducting processor. The qubits are effectively coupled to each other through virtual photon exchange, mediated by a bus resonator dispersively coupled to the qubits. By tuning the competition between the external continuous driving and intra-qubit interaction, we can move the system from the $Z_{2}$-symmetric phase to the symmetry-breaking phase, where the two degenerate highest eigenstates of the governing Hamiltonian, superimposed with each other, form a GHZ-like cat state. We characterize the SSB extent with the average two-qubit longitudinal correlation. The quantum nature of the SSB phase is confirmed by the measured multi-qubit quantum correlation, as well as by the reconstructed multi-qubit Wigner function, which exhibits interference fringes with alternating positive and negative values. Besides fundamental interest, the demonstrated SSB process can be used for noise-insensitive many-body entanglement manipulation.